# First Measurement of Entanglement Dynamics in the SYK Model using Quantum Computers

### v2

@@ Abstract
The Sachdev-Ye-Kitaev (SYK) model, fundamental to quantum chaos, many-body physics, and holographic duality, exhibits both maximal chaos and exact solvability, yet its entanglement entropy has never been measured experimentally. Here, we present the first measurement of entanglement entropy growth under the SYK Hamiltonian using IBM’s superconducting quantum computer as a programmable quantum platform. We implement an optimized swap-based many-body interference protocol tailored for limited qubit connectivity to measure entanglement entropy. Additionally, we enhance the randomized measurement protocol for entropy measurement by expanding the unitary ensemble, which increases the circuit volume and pushes the capability of the near-term devices. However, we tackle this challenge by employing a quantum multi-programming strategy that parallelizes circuit execution. This work paves the way for exploring chaotic quantum dynamics and establishes scalable methods for investigating complex entanglement growth on current quantum platforms.
Introduction.— Quantum entanglement [1] plays a crucial role in various fields, including quantum computing, quantum information science, many-body physics, high-energy and nuclear physics, and black hole physics. Additionally, effectively harnessing quantum information and entanglement dynamics is essential for advancing quantum technologies and exploring complex quantum systems [2, 3]. However, the exponential growth of the Hilbert space associated with increasing system size poses significant challenges for classical computers, making it difficult to compute entanglement entropy for highly entangled states. Besides, experimentally measuring the entanglement entropy and its dynamics in complex quantum systems is challenging due to the specific requirements for state preparation, the execution of time evolution, and the implementation of entropy measurement techniques, such as those based on quantum gas microscopy [4, 5].
@@
In this regard, quantum computers can be utilized as experimental probes to measure the entanglement entropy of the SYK model. In fact, real quantum devices are emerging as highly effective programmable quantum platforms, capable of accurately preparing complex quantum many-body models, and offer a valuable testing ground for novel phenomena that would be difficult or impossible to realize through experimental means [13, 14, 15]. Programmable quantum platforms are being used to study discrete time crystals [16, 17], many-body localization [18, 19, 20, 21], quantum phases, and topologically ordered states [22, 23, 24, 25, 26, 27, 28] in quantum many-body systems. Building on our previous work that investigated the Page Curve entanglement dynamics using quantum computers as programmable platform [29], we now shift our focus to a more complex measurement of entanglement dynamics in the SYK model. Although some studies [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have addressed the SYK model, we are the first to measure the entanglement entropy within this model using a quantum computer. Specifically, we report our findings on the growth of entanglement entropy for subsystems of varying sizes over time, utilizing IBM’s superconducting quantum computer. This measurement is performed while evolving a total product state with the time evolution operator associated with the SYK Hamiltonian.
@@
The Model.— The SYK model is given by the Hamiltonian,
@@
$H=\sum_{n=1}^{M}a_{n}\hat{P}_{n},$ (3)
where $M$ is the number of terms in Eq. (1), $\hat{P}_{n}$ is the Pauli operator of weight $N/2$, i.e. it acts on the Hilbert space of $N/2$ qubits, and $a_{n}$ is the corresponding coefficient.
Here, we introduce the entanglement entropy measure we use in this study. Consider a system of $N$ qubits in a pure state $|\psi\rangle$, with its density matrix given by $\rho=|\psi\rangle\langle\psi|$. If we partition this system into two subsystems consisting of $L$ and $M=N-L$ qubits, the entanglement entropy, specifically the $n$-Rényi entropy associated with the $L$-qubit subsystem, is defined as
@@
where $\rho_{L}$ is the reduced density matrix for the subsystem with $L$ qubits. This reduced density matrix is obtained by tracing out the subsystem of $M=N-L$ qubits, such that $\rho_{L}=\mathrm{Tr}_{M}\rho$. Our focus in this study is $n=2$, corresponding to the 2-Rényi entropy (commonly referred to as Rényi entropy), $S^{(2)}$. Additionally, as $n$ approaches 1, the $n$-Rényi entropy converges to the von Neumann entropy:
@@
$U(t)=e^{-iHt}\approx\left(\prod_{n=1}^{M}e^{-iH_{n}t/r}\right)^{r},$ (6)
@@
Entanglement Entropy Measuring Protocols.— We consider two protocols to measure the purity and Rényi entropy of subsystems in a total quantum state: 1. swap-based many-body interference (SWAP-MBI) protocol and 2. randomized measurement (RM) protocol.
@@
IBMQ Results.— Despite the importance of the SYK model and the entanglement entropy measurement, implementing them on near-term quantum computers is challenging due to the complexity of the Pauli strings and the computation overhead of the two entanglement entropy measurement protocols, the SWAP-MBI and the RM protocols. In particular, one Trotter step has 126 circuit depths and 39 two-qubit gates (CX or CZ). Since we compute up to five Trotter steps, the largest circuit depth and the number of two-qubit gates are 630 and 195, respectively. Based on the Trotter step implementation, the SWAP-MBI protocol needs 196 additional circuit depth to implement the SWAP test in the protocol. On the other hand, the RM protocol needs only 5 additional circuit depths. However, the protocol needs a large number of repetitions for the random sampling of the $SU(2)$ circuits. To overcome those challenges, we apply several methods detailed below.
We conduct our experiment on the IBM Quantum computer, ibm_marrakesh which has 156 qubits with basis gates CZ, ID, RX, RZ, RZZ, SX, X. We evolve a product state of $N_{\mathrm{qubit}}=3$ that corresponds to $N=6$ Majorana fermions, with time-evolution operator driven by the SYK Hamiltonian via the first-order Trotterization up to $t=10$ (in unit of $\hbar/E$ where $E$ is the overall energy scale of the SYK Hamiltonian) in five Trotter steps (Trotter step size 2.0). In each Trotter step, we employ the entanglement entropy measuring protocols: SWAP-MBI and RM protocols to measure the entropy of left single-qubit and two-qubit subsystems, respectively. As going beyond the case $N>6$ for the SYK model is highly challenging on near-term quantum computers due to the polynomial growth of gate complexity with $N$, we focus on the feasible case of $N=6$ Majorana fermions or $N_{\mathrm{qubit}}=3$ qubits. Additionally, we employed an optimized version of the SWAP-MBI protocol, designed specifically for quantum devices with limited qubit connectivity [29] as shown in Fig. 3.
{quantikz}
[…]
@@
Still, running quantum algorithms on near-term quantum devices, such as IBM Quantum processors, poses a significant challenge due to the errors and noise present in these systems. But, we precisely measure the variation of entanglement entropy with time under the SYK Hamiltonian evolution in quantum computers up to long circuit depths (more than 600 depths, five timesteps) by applying four quantum error mitigation methods to mitigate the errors and noise inherent to quantum devices: Zero-Noise Extrapolation (ZNE), Pauli Twirling (PT), Dynamic Decoupling (DD), and Matrix-free Measurement Mitigation (M3). Refer to Ref. [43] for further details.
@@
During our measurement of the subsystem entanglement entropy over time, we discovered that the SWAP-MBI protocol is more accurate for measuring entropy at early times when it is relatively small. In contrast, the RM protocol, which relies on local Haar-random unitaries, tends to produce incorrect results due to significant statistical fluctuations in its purity estimation during these early times.
As illustrated in Fig. 4, there is excellent agreement between the exact and measured Rényi entropy $S^{(2)}$ on the ibm__\__marrakesh device. The slight deviation observed at larger times (specifically, $t=6-10$) can be attributed to inherent Trotter error. However, at these times, the noiseless Qiskit simulation and the error-mitigated measured values of the SWAP-MBI protocol align very well.
@@
We measure the Rényi entropy using the RM protocol for $t=6,8,$ and $10$, as shown in Fig. 5. Again, we see excellent agreement between exact results and results obtained by RM protocol at ibm__\__marrakesh. A key challenge of the RM protocol is its reliance on an ensemble of local Haar-random unitaries to minimize the variance in entropy computation. In our experiments, we utilize $150$ samples of the RM protocol circuit. Besides, as we apply the Zero Noise Extrapolation (ZNE) and the Pauli Twirling methods to mitigate errors, each method replicates the variation three times (ZNE) and ten times (PT), respectively. Therefore, the total number of circuits for the RM protocol at each measurement time is 4,500 ($150\times 3\times 10$). Since we have three measurement points (6, 8, and 10), the total number of circuits is 135,000.
@@
Conclusion and Outlook.— Using IBM’s superconducting quantum computer, we have achieved the first experimental measurement of entanglement entropy growth in the SYK model, which marks a significant milestone in the quantum simulation of chaotic many-body systems. Despite the challenges posed by the dense interaction structure of the SYK Hamiltonian and the computational demands of entanglement entropy protocols, we successfully implemented both the SWAP-based many-body interference (SWAP-MBI) and randomized measurement (RM) protocols on IBM’s superconducting quantum hardware. To address the limitations of qubit connectivity, we employed an optimized version of the SWAP-MBI protocol and extracted meaningful entanglement dynamics from quantum circuits of large circuit depths. Importantly, we tackled the substantial resource demands of the RM protocol by utilizing Quantum Multi-Programming (QMP) to parallelize 135,000 circuits. This approach significantly enhanced the efficiency of entropy estimation without compromising accuracy. Our results show excellent agreement with noiseless simulations, validating the SWAP-MBI and RM protocols across different entanglement regimes and timescales.

### v3

# Probing Entanglement Dynamics in the SYK Model using Quantum Computers

@@ Abstract
Quantum computers are expected to be vital for exploring complex dynamics in many-body quantum systems. Thus, validating established results on current quantum computers is essential for evaluating their future utility. Hence, we investigate the entanglement entropy of the Sachdev-Ye-Kitaev (SYK) model, a paradigmatic model of quantum chaos, many-body physics, and holographic duality, in current IBM’s superconducting quantum computers. We implement optimized swap-based many-body interference protocol and randomized measurement protocol tailored for IBM quantum computers’ limited qubit connectivity. Additionally, we employ quantum multi-programming that parallelizes circuit execution to improve the results obtained by the randomized measurement protocol. Finally, by incorporating the quantum error mitigation techniques into our implementation of the entropy measurement protocols on IBM quantum hardware, we show that the current noisy quantum computer can yield results aligned with theoretical expectations, therefore affirming its capability to explore chaotic quantum dynamics in complex quantum systems.
journal: Results in Physics
1 Introduction
Quantum entanglement [1] plays a crucial role in various fields, including quantum computing, quantum information science, many-body physics, high-energy and nuclear physics, and black hole physics. Additionally, effectively harnessing quantum information and entanglement dynamics is essential for advancing quantum technologies and exploring complex quantum systems [2, 3]. However, the exponential growth of the Hilbert space associated with increasing system size poses significant challenges for classical computers, making it difficult to compute entanglement entropy for highly entangled states. Besides, experimentally measuring the entanglement entropy and its dynamics in complex quantum systems is challenging due to the specific requirements for state preparation, the execution of time evolution, and the implementation of entropy measurement techniques, such as those based on quantum gas microscopy [4, 5].
@@ 1 Introduction
In this regard, quantum computers can be utilized as experimental probes to measure the entanglement entropy of the SYK model. In fact, real quantum devices are emerging as highly effective programmable quantum platforms, capable of accurately preparing complex quantum many-body models, and offer a valuable testing ground for novel phenomena that would be difficult or impossible to realize through experimental means [13, 14, 15]. Programmable quantum platforms are being used to study discrete time crystals [16, 17], many-body localization [18, 19, 20, 21], quantum phases, and topologically ordered states [22, 23, 24, 25, 26, 27, 28] in quantum many-body systems. Building on our previous work that investigated the feasibility of current quantum computers as programmable platform to study the Page Curve entanglement dyanicms [29], we now shift our focus to a more complex measurement of entanglement dynamics in the SYK model. Although some studies [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have addressed the time-evolution of the SYK model in quantum computers, we employ an optimized version of swap-based measurement protocol, delineated in [29], and randomized measurement protocols  [40, 41, 42, 43, 44, 45, 46] to probe the the entanglement entropy within this model using IBM’s superconducting quantum computer. Specifically, we study the growth of entanglement entropy for subsystems of varying sizes over time by evolving a total product state using the time evolution operator associated with the SYK Hamiltonian.
@@
The structure of the paper is as follows. In section 2, we provide a telegraphic description of the SYK model and the entanglement entropy measure, namely the second Rényi entropy that we compute in IBM’s superconducting quantum computers. Section 3 describes the two entanglement entropy measuring protocols adopted in this work. In section 4, we present our results of the variation of the entanglement entropy with time under the time evolution by the SYK Hamiltonian, which we have obtained by employing the entanglement measuring protocols in combination with quantum error mitigation on IBM’s quantum devices. Finally, we conclude in section 5.
2 The SYK Model
The SYK model is given by the Hamiltonian,
@@ 2 The SYK Model
$H=\sum_{n=1}^{Q}a_{n}\hat{P}_{n},$ (3)
where $Q$ is the number of terms in Eq. (1), $\hat{P}_{n}$ is the Pauli operator of weight $N/2$, i.e. it acts on the Hilbert space of $N/2$ qubits, and $a_{n}$ is the corresponding coefficient.
Here, we introduce the entanglement entropy measure we use in this study. Consider a system of $N$ qubits in a pure state $|\psi\rangle$, with its density matrix given by $\rho=|\psi\rangle\langle\psi|$. If we partition this system into two subsystems consisting of $L$ and M=NLM=N-L qubits, the entanglement entropy, specifically the $n$-Rényi entropy associated with the $L$-qubit subsystem, is defined as
@@
where $\rho_{L}$ is the reduced density matrix for the subsystem with $L$ qubits. This reduced density matrix is obtained by tracing out the subsystem of M=NLM=N-L qubits, such that $\rho_{L}=\mathrm{Tr}_{M}\rho$. Our focus in this study is $n=2$, corresponding to the 2-Rényi entropy (commonly referred to as Rényi entropy), $S^{(2)}$. Additionally, as $n$ approaches 1, the $n$-Rényi entropy converges to the von Neumann entropy:
@@
$U(t)=e^{-iHt}\approx\left(\prod_{n=1}^{Q}e^{-iH_{n}t/r}\right)^{r},$ (6)
@@
3 Entanglement Entropy Measuring Protocols
We consider two protocols to measure the purity and Rényi entropy of subsystems in a total quantum state: 1. swap-based many-body interference (SWAP-MBI) protocol and 2. randomized measurement (RM) protocol.
@@ 3 Entanglement Entropy Measuring Protocols
4 IBMQ Results
Despite the importance of the SYK model and the entanglement entropy measurement, implementing them on near-term quantum computers is challenging due to the complexity of the Pauli strings and the computation overhead of the two entanglement entropy measurement protocols, the SWAP-MBI and the RM protocols. In particular, one Trotter step has 126 circuit depths and 39 two-qubit gates (CX or CZ). Since we compute up to five Trotter steps, the largest circuit depth and the number of two-qubit gates are 630 and 195, respectively. Based on the Trotter step implementation, the SWAP-MBI protocol needs 196 additional circuit depth to implement the SWAP test in the protocol.
On the other hand, the RM protocol needs only 5 additional circuit depths. However, the protocol needs a large number of repetitions for the random sampling of the $SU(2)$ circuits. To overcome those challenges, we apply several methods detailed below.
We conduct our experiment on the IBM Quantum computer, ibm_marrakesh which has 156 qubits with basis gates CZ, ID, RX, RZ, RZZ, SX, X. We evolve a product state of $N_{\mathrm{qubit}}=3$ that corresponds to $N=6$ Majorana fermions, with time-evolution operator driven by the SYK Hamiltonian via the first-order Trotterization up to $t=10$ (in unit of $\hbar/E$ where $E$ is the overall energy scale of the SYK Hamiltonian) in five Trotter steps (Trotter step size 2.0). We also present the quantum circuit associated with the single Trotter step in Fig. 3. In each Trotter step, we employ the entanglement entropy measuring protocols: SWAP-MBI and RM protocols to measure the entropy of left single-qubit and two-qubit subsystems, respectively. As going beyond the case $N>6$ for the SYK model is highly challenging on near-term quantum computers due to the polynomial growth of gate complexity with $N$, we focus on the feasible case of $N=6$ Majorana fermions or $N_{\mathrm{qubit}}=3$ qubits. Additionally, we employed an optimized version of the SWAP-MBI protocol, designed specifically for quantum devices with limited qubit connectivity [29] as shown in Fig. 4.
Figure 3: The quantum circuit associated with the single Trotter step for $N=6$ Majorana fermions ($N_{\mathrm{qubit}}=3$) for a single instance of SYK Hamiltonian. The circuit diagram is generated by qiskit.visualization.circuit_\_drawer. The circuit is continued from the upper panel to the lower panel.
@@ 4 IBMQ Results
Still, running quantum algorithms on near-term quantum devices, such as IBM Quantum processors, poses a significant challenge due to the errors and noise present in these systems. However, we precisely measure the variation of entanglement entropy with time under the SYK Hamiltonian evolution in quantum computers up to long circuit depths (more than 600 depths, five timesteps) by applying four quantum error mitigation methods to mitigate the errors and noise inherent to quantum devices: Zero-Noise Extrapolation (ZNE), Pauli Twirling (PT), Dynamic Decoupling (DD), and Matrix-free Measurement Mitigation (M3). Additionally, we conduct a comprehensive performance comparison of ZNE, PT, and DD for the Hamiltonian simulation on IBM quantum computers, as discussed in [60], where we show that combining these Quantum Error Mitigation (QEM) methods, i.e., ZNE, PT, and DD, yields the highest accuracy. For more details on the QEM implementations, please refer to Refs. [50, 60].
@@
During our measurement of the subsystem entanglement entropy over time, we discovered that the SWAP-MBI protocol is more accurate for measuring entropy at early times when it is relatively small. In contrast, the RM protocol, which relies on local Haar-random unitaries, tends to produce incorrect results due to significant statistical fluctuations in its purity estimation during these early times because when the time $t$ is small the quantum state is close to the initial product state, so the Rényi entropy associated with its subsystem is relatively small (i.e. the purity is close to 1), and such small values of the Rényi entropy are prone to larger statistical errors as pointed out in [41, 44].
As illustrated in Fig. 5, there is excellent agreement between the exact and measured Rényi entropy $S^{(2)}$ on the ibm_\_marrakesh device. The slight deviation observed at larger times (specifically, t=610t=6-10) can be attributed to inherent Trotter error. However, at these times, the noiseless Qiskit simulation and the error-mitigated measured values of the SWAP-MBI protocol align very well.
@@
We measure the Rényi entropy using the RM protocol for $t=6,8,$ and $10$, as shown in Fig. 6. Again, we see excellent agreement between exact results and results obtained by RM protocol at ibm_\_marrakesh. A key challenge of the RM protocol is its reliance on an ensemble of local Haar-random unitaries to minimize the variance in entropy computation. In our experiments, we utilize $150$ samples of the RM protocol circuit. Besides, as we apply the Zero Noise Extrapolation (ZNE) and the Pauli Twirling (PT) methods to mitigate errors, each method replicates the variation three times (ZNE) and ten times (PT), respectively. Therefore, the total number of circuits for the RM protocol at each measurement time is 4,500 ($150\times 3\times 10$). Since we have three measurement points (6, 8, and 10), the total number of circuits is 135,000.
@@
5 Conclusion and Outlook
Using IBM’s superconducting quantum computer, we have obtained the measurement of entanglement entropy growth in the SYK model, which are in excellent agreement with the theoretical results. Despite the challenges posed by the dense interaction structure of the SYK Hamiltonian and the computational demands of entanglement entropy protocols due to the current hardware limitations, we successfully implemented both the SWAP-based many-body interference (SWAP-MBI) and randomized measurement (RM) protocols on IBM’s superconducting quantum hardware. To address the limitations of qubit connectivity, we employed an optimized version of the SWAP-MBI protocol and extracted meaningful entanglement dynamics from quantum circuits of large circuit depths. Importantly, we tackled the substantial resource demands of the RM protocol by utilizing Quantum Multi-Programming (QMP) to parallelize 135,000 circuits. This approach significantly enhanced the efficiency of entropy estimation without compromising accuracy. Finally, our results show excellent agreement with noiseless simulations, validating the SWAP-MBI and RM protocols across different entanglement regimes and timescales. Moreover, the von Neumann entropy in the SYK model can be computed using superconducting quantum computers following [64, 65].